# BubbleONet: A Physics-Informed Neural Operator for High-Frequency Bubble Dynamics

### v1

@@ 6 Conclusions
Bubble dynamics plays a critical role in many scientific and engineering domains, where rapid and accurate modeling is essential for understanding phase interactions and energy transfer in multiphase systems. While well-characterized by mathematical models such as the Rayleigh–Plesset and Keller–Miksis equations, the lack of efficient prediction methods for multi-bubble simulations limits its applicability in real-time and in-situ scenarios. In this paper, we propose a physics-informed neural operator model, BubbleONet, which leverages the function approximation capabilities of DeepONet to model macroscopic bubble dynamics. The framework accommodates bubbles governed by both the Rayleigh-Plesset equation and Keller-Miksis equation, and handles scenarios involving single and multiple initial radii. Key findings include:
@@
    The implementation of the Rowdy activation function effectively addresses spectral bias, enabling the model to accurately predict bubble dynamics across a wide range of pressure amplitudes and frequencies.
@@
    A comparison between single-step and two-step training techniques for BubbleONet demonstrates that the two-step approach significantly enhances the training process. It achieves faster convergence while maintaining comparable accuracy across validation, interpolation, and extrapolation scenarios.
@@
    The model is more sensitive to the frequency of the pressure than to its amplitude. Both training methods can capture natural and driving frequencies when extrapolating over amplitude; however, they struggle to predict accurately when extrapolating over frequency.
Despite its promising performance, the model exhibits several limitations that warrant further investigation. Notably, it currently fails to accurately capture bubble dynamics excited at its resonance frequency, which limits its current applicability in general turbulent flows, where a broad band spectrum of exciting frequencies, including the resonant frequency, is typically present. Further, its predictive capability remains constrained to the time domain represented in the training data, resulting in significant errors when evaluated on out-of-distribution samples. Ongoing research aims to address these challenges by improving the model’s adaptability and generalization to unseen scenarios.

### v3

@@ 6 Conclusions
High-frequency bubble dynamics plays an important role in many scientific and engineering domains, where rapid and accurate modeling is essential for understanding phase interactions and energy transfer in multiphase systems. While mathematical models such as the Rayleigh–Plesset and Keller–Miksis equations provide a well-characterized description of bubble dynamics, the lack of efficient prediction methods for multi-bubble simulations limits their applicability in real-time and in-situ scenarios. In this work, we employed a physics-informed two-step DeepONet model, which leverages the function approximation capabilities of DeepONet to model bubble dynamics. The framework accommodates bubbles governed by both the Rayleigh-Plesset and Keller-Miksis equations and handles scenarios involving single and multiple initial radii. In particular, we solved bubble dynamics over a large range of amplitudes ($1,10]\times 10^{5}$ Pa) and frequencies ($200,2000]\times 10^{6}$ KHz), representing a wide range of physically realistic scenarios. Our key findings show that the Rowdy activation function effectively mitigates spectral bias, enabling accurate prediction of bubble dynamics across varying pressure amplitudes and frequencies. The two-step training approach improves efficiency, achieving faster convergence while maintaining comparable accuracy in validation, interpolation, and extrapolation. The model is more sensitive to driving frequency than amplitude, capturing natural and driving frequencies when extrapolating over amplitude but facing challenges with frequency extrapolation. While these findings hold for the two-step DeepONet architecture, several important insights emerge for the proposed two-step DeepOKAN framework:
@@
    The proposed two-step DeepOKAN framework leverages the KAN architecture, providing parametric efficiency and interpretability absent in equivalent MLP-based models. However, the KAN architecture still exhibits spectral bias similar to MLPs.
@@
    To address high-frequency spectral bias, we introduce a strategy that continually trains the trunk network, improving generalization across a range of frequencies and pressure amplitudes. This procedure can be extended to other problems without loss of generality using the KAN architecture.
@@
    Although the RBF basis [abueidda2025deepokan] is computationally efficient, it is limited in learning high-frequency phenomena such as bubble dynamics. We therefore employ a spline basis in the trunk network for improved training performance, while RBF is used in the branch network. In the two-step framework, the trunk network constructs the basis to represent high-frequency dynamics accurately, and the branch network weights this basis according to the input pressure distribution. The branch network’s quality is dependent on the trunk network, given the sequential training procedure.
Despite its promising performance, the proposed two-step DeepOKAN model exhibits limitations. Importantly, it fails to capture bubble dynamics at resonance frequency, restricting its applicability in turbulent flows where a broad spectrum of frequencies (including resonance) is present. Furthermore, predictive capability is constrained to the time domain represented in the training data, leading to significant errors on out-of-distribution samples. Future research aims to enhance the model’s adaptability and generalization to unseen scenarios.
Acknowledgement
We would like to thank all the reviewers for their valuable feedback and constructive suggestions, which have helped improve the quality and clarity of this work.